How much faster does light travel in a crystal with refraction index 1.48 than in another with refraction index 2.29?. Please include units
Answer #1

since u=c/v

uv=constant

u1v1=u2v2

v2=u1v1/u2

=1.48*v1/2.29

v2 =0.6462 * v1

that in second medium light will travell 0.6462 time faster than the first medium.
Answer #2

Refractive index (u) = ( speed of light in vacuum/air (c)) / (speed of light in medium ( V)

Let u1 = refractive index of crystal. = 1.48 ,                  c = speed of light in vaccum/air = 3*108 m/s

u1 = c / V1                          ( v1 = velocity of light in crystal)

V1 = c / u1 = 3*108 m/s / 1.48

V1 = 2.02*108 m/s

u2 = refractive index if other material = 2.29

V2 = velocity of light in other material

u2 = c/V2

V2 = c /u2 = 3*108 m/s / 2.29

V2 = 1.310*108 m/s

V1 /V2 = 2.02*108 / 1.310*108 = 1.541

V1 = 1.541 V2

Hence velocity of light in crystal is 1.541 times greater than velocity in other material.